OpenStudy (anonymous):

MEDAL!!!!!!! 1. Solve for x. 4(x – b) = x (1 point) b = 4/3 x b = 3/4 x x =4/3 b x = 3/4 b 2. Solve for y, then find the value of y when given x = –2. 6x = 7 – 4y (1 point) –12 7/4 19/4 24 3. Solve for h. A = 1/2bh (1 point) h= 2a/b 2A =b/h h = 1/2 Ab h = 2Ab 4. Find the width of a rectangle with a perimeter of 90 and a length of 15. (1 point) 90 15 45 30

OpenStudy (anonymous):

even if you just know one of the answers, please reply!!!

OpenStudy (ajprincess):

4(x – b) = x Using distributive property distribute the expression on the left side Then u will get 4*x-4*b=x 4x-4b=x To solve for x u have to isolate x. to do that follow the steps 1.Add 4b both sides 2.Subtract x both sides 3.Divide by 3 both sides Can u do it @kittykatchan

OpenStudy (anonymous):

well it's worth a shot. hold on a sec and lemme try :)

OpenStudy (ajprincess):

sure:)

OpenStudy (anonymous):

umm, is it x = 3/4 b

OpenStudy (nubeer):

yes

OpenStudy (anonymous):

are you sure, nubeer?

OpenStudy (nubeer):

ohh lol no..sorry my bad.

OpenStudy (anonymous):

okay, so then it's not that answer?

OpenStudy (ajprincess):

nope

OpenStudy (anonymous):

fine, then is it x = 4/3 b???

OpenStudy (ajprincess):

yup it is:) 4x-4b=x 4x-4b+4b=x+4b 4x=x+4b 4x-x=x+4b-x 3x=4b x=4b/3

OpenStudy (anonymous):

oh, okay great. thanks! :) can you help with any others or do you know only that one?

OpenStudy (ajprincess):

for the second one plug in the given value in place of x in 6x = 7 – 4y and solve for y. 6*(-2)=7-4y Can u find y @kittykatchan?

OpenStudy (anonymous):

yeah, probably. i was just figuring out the last one. one sec and lemme try

OpenStudy (anonymous):

okay, i'll probably have to look in my book for that one.

OpenStudy (anonymous):

it's 19/4 right???

OpenStudy (ajprincess):

ya right:) 6*(-2)=7-4y -12=7-4y -12-7=7-4y-7 -19=-4y -19/-4=-4y/-4 19/4=y

OpenStudy (anonymous):

woo hoo! okay, all i have left is the third one. do you know it?? :)

OpenStudy (ajprincess):

so u have finishd the 4th one? good:)

OpenStudy (anonymous):

yup, i already finished it. :) if you don't know the third one, it's ok i'll just guess.

OpenStudy (anonymous):

nubeer, why are you creeping? lol

OpenStudy (ajprincess):

well give it a try if u find t difficult plz ask then I will help u:)

OpenStudy (anonymous):

it's ok, i just gueesed. thanks for your help ajprincess! heres your medal :D